Please help with the homework:
Peter get employed on the first of January 2020, suppose Peter purchase an apartment of R180000 with a deposit of R45000. Peter secure a mortgage bond with a bank for the balance at 18% interest per annum, compounded monthly with a term Of 20 years. What will be the monthly payment?

Answer by math_tutor2020(3817)


Peter made a down payment of R45000
We'll subtract this from the cost of the apartment.

R180000 - R45000 = R135000

It might be helpful to think of that above calculation as
180 - 45 = 135
then just stick "thousand" on to each value

The value R135000 represents the amount Peter borrows.

L = 135,000 = loan amount
P = unknown monthly payment
r = annual interest rate in decimal form = 0.18
i = monthly interest rate in decimal form = r/12 = 0.18/12 = 0.015 which is exact
n = number of months = 20*12 = 240 (equivalent to 20 years)

Let's calculate the value of P
P = (L*i)/(1 - (1+i)^(-n))
P = (135,000*0.015)/(1 - (1+0.015)^(-240))
P = 2,083.47055654027
P = 2,083.47

Answer: The monthly payment is R2083.47